OpenStudy (anonymous):

solve by completing the square x^2+(3/2)x=3

OpenStudy (anonymous):

really?? ok then, this is not the easiest way to get this answer but x^2+(3/2)x-3=0 (x+3/4)^2-9/16-3=0 (x+3/4)^2=57/16 \[x+3/4=\pm \sqrt{57/16}\] \[x=-3/4 \pm \sqrt{57}/4\]

OpenStudy (anonymous):

x=1.14 or x=-2.64